OpenStudy (anonymous):

x – 3y = 10 5x – 3y = 14

OpenStudy (anonymous):

i have to use the addition method?

OpenStudy (anonymous):

are u using mymathlab.com ?

OpenStudy (anonymous):

it would be 6x=24

OpenStudy (anonymous):

and then divide by 6

OpenStudy (anonymous):

which is 4

OpenStudy (anonymous):

subtract the second equation from the first and you'll get -4x = -4 x=1 then substitute 1-3y=10 -3y=9 y=-3

OpenStudy (anonymous):

ok i see how that worked that you...so it like the distribution process?